# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

### 03 Mar 2026 v2

@@ Appendix C Quadratic $L$-Function Zeros to 20 Decimal Places (Seven Discriminants)
11 $4.89997399700703650104$ $1\times 10^{-31} 1111$ $26.95853518038046741969$ $8.98\times 10^{-32}$
22 $7.62842884176939783416$ $1.32\times 10^{-31}$ $12$ $28.09744496063079735462$ $2.57\times 10^{-31}$
33 $10.8065881638617120144$ $1.52\times 10^{-31}$ $13$ $29.93076421015190503723$ $2.66\times 10^{-31}$
44 $12.3105429942365296811$ $1.31\times 10^{-31}$ $14$ $31.63813949132101731704$ $2.90\times 10^{-32}$
55 $15.1957542506451227685$ $2.61\times 10^{-31}$ $15$ $33.84563089515844186412$ $7.91\times 10^{-32}$
66 $17.0222859743083473390$ $1.70\times 10^{-31}$ $16$ $34.74577700617002330947$ $5.29\times 10^{-32}$
77 $18.8059589077071483998$ $4.50\times 10^{-32}$ $17$ $36.54166385177458968906$ $2.68\times 10^{-31}$
88 $21.1316459622213438826$ $2.91\times 10^{-31}$ $18$ $38.77557777387006581424$ $1.96\times 10^{-31}$
99 $23.0838499962005465429$ $1.06\times 10^{-31}$ 1919 $39.78688099766664071794$ $2.48\times 10^{-31}$
$10$ $24.2019635578156016126$ $1.91\times 10^{-31}$ 2020 $41.34220616442193704510$ $3.51\times 10^{-31}$
@@
11 $3.80462763305086509714$ $5.15\times 10^{-32}$ $11$ $23.56131971313784471635$ $3.18\times 10^{-31}$
22 $6.69222332050013115991$ $2\times 10^{-31} 1212$ $25.41163389239269596203$ $2.13\times 10^{-31}$
33 $8.89059295872674150934$ $2.54\times 10^{-31}$ $13$ $27.01394398585067234435$ $2.92\times 10^{-31}$
44 $11.1883927450749091542$ $2.58\times 10^{-31}$ $14$ $28.44220325774560811851$ $2.64\times 10^{-31}$
55 $12.9661788080288453750$ $2.81\times 10^{-31}$ $15$ $30.20400655643888185431$ $2.43\times 10^{-31}$
66 $15.1814808758882167354$ $1.02\times 10^{-31}$ $16$ $31.64807761494068636690$ $5.56\times 10^{-33}$
77 $16.6326332745237621232$ $5.47\times 10^{-32}$ $17$ $33.03713287950650360192$ $1.05\times 10^{-31}$
88 $18.8843694571206508082$ $3.57\times 10^{-32}$ $18$ $35.02737848501270258165$ $6.67\times 10^{-32}$
99 $20.1039281912458185731$ $9.81\times 10^{-32}$ 1919 $35.77804457652351027242$ $2.50\times 10^{-32}$
$10$ $22.2858391072268876587$ $1.51\times 10^{-31}$ 2020 $37.92681682137521283529$ $3.25\times 10^{-31}$
@@
11 $3.11934147900860341390$ $4.45\times 10^{-32}$ $11$ $23.59202788517293715039$ $3.77\times 10^{-32}$
22 $7.23159073941876201503$ $1.17\times 10^{-31}$ $12$ $25.37170440577142621541$ $4.33\times 10^{-32}$
33 $8.62542663503259159735$ $4.11\times 10^{-32}$ $13$ $26.38431352575075899279$ $6.00\times 10^{-32}$
44 $10.3364207262315390298$ $1.36\times 10^{-31}$ $14$ $27.64980842472696812713$ $2.33\times 10^{-31}$
55 $12.6170127910231787304$ $4.72\times 10^{-32}$ $15$ $29.39145603391187839556$ $3.70\times 10^{-31}$
66 $15.1483324170057442246$ $2.53\times 10^{-31}$ $16$ $31.01874209224480391871$ $8.86\times 10^{-32}$
77 $16.2748260574985884179$ $8.23\times 10^{-32}$ $17$ $32.42341728496373366098$ $2.06\times 10^{-31}$
88 $18.7512523562342329165$ $1.65\times 10^{-32}$ $18$ $34.46644233774365394409$ $4.03\times 10^{-31}$
99 $19.5480414433470398699$ $1.19\times 10^{-31}$ 1919 $35.76319303421611596021$ $6.16\times 10^{-32}$
$10$ $20.9591819174050311814$ $3.35\times 10^{-31}$ 2020 $36.72754705153707320059$ $1.42\times 10^{-31}$
@@
11 $2.68865813246759758667$ $41\times 10^{-31} 11$11 $20.03461157675018829473$ $1.62\times 10^{-31}$
22 $5.29243117677719815160$ $7\times 10^{-31} 1212$ $21.55547824218793740140$ $3.62\times 10^{-31}$
33 $6.97192424379172752833$ $1.20\times 10^{-31}$ $13$ $22.66785732205079621070$ $9.70\times 10^{-32}$
44 $9.22463743993556115082$ $2\times 10^{-31} 1414$ $24.40914964223159369360$ $2.18\times 10^{-31}$
55 $10.4457215184129248230$ $2.97\times 10^{-31}$ $15$ $25.67129733212202220561$ $2.53\times 10^{-31}$
66 $12.6311537181665105855$ $3.28\times 10^{-31}$ $16$ $26.75610741237823107924$ $3.22\times 10^{-31}$
77 $13.7771527057734958652$ $2.51\times 10^{-31}$ $17$ $28.86856257427073136901$ $1.56\times 10^{-31}$
88 $15.6036569703702537829$ $3.03\times 10^{-31}$ $18$ $29.21969083370577778290$ $1.10\times 10^{-31}$
99 $17.0989880869194034869$ $1.05\times 10^{-31}$ 1919 $31.08725180031803397862$ $4.04\times 10^{-31}$
$10$ $18.4632988501214004408$ $1.83\times 10^{-31}$ 2020 $32.42353686755135030794$ $3.61\times 10^{-31}$
@@
11 $2.77728324207659233094$ $6.55\times 10^{-33}$ $11$ $19.53861609989941145163$ $2.70\times 10^{-31}$
22 $4.52677408701547610929$ $2.86\times 10^{-31}$ $12$ $20.39328285450433247153$ $7.16\times 10^{-32}$
33 $7.29088506700227937479$ $2.61\times 10^{-31}$ $13$ $21.98889814760568313347$ $3.40\times 10^{-31}$
44 $8.39512407283755124263$ $1.74\times 10^{-31}$ $14$ $23.69153462655719640538$ $2.67\times 10^{-32}$
55 $10.1163068547312977911$ $2.08\times 10^{-31}$ $15$ $24.73886680489813285047$ $3.18\times 10^{-31}$
66 $12.0220620381859538014$ $3.73\times 10^{-32}$ $16$ $26.38179066809542195017$ $1.65\times 10^{-31}$
77 $13.5442863718297874149$ $5.68\times 10^{-32}$ $17$ $27.02957442259405572955$ $1.67\times 10^{-31}$
88 $15.0682949619190434065$ $1.90\times 10^{-31}$ $18$ $28.72073200609003118776$ $5.74\times 10^{-31}$
99 $16.0162614771505928089$ $9\times 10^{-32} 1919$ $30.39449098988413064026$ $3.27\times 10^{-31}$
$10$ $18.2004252754926244593$ $1.75\times 10^{-31}$ 2020 $31.36586297616383901441$ $2.58\times 10^{-31}$
@@
11 $2.48821020930540853564$ $9.61\times 10^{-32}$ $11$ $18.18218004680212107005$ $2.46\times 10^{-32}$
22 $4.03159701418841483371$ $3.83\times 10^{-32}$ $12$ $19.28135808746821972149$ $3.45\times 10^{-31}$
33 $6.33545319091356197520$ $3.06\times 10^{-31}$ $13$ $20.40160148601184344650$ $3.27\times 10^{-31}$
44 $7.98779634495761720535$ $1.35\times 10^{-31}$ $14$ $21.56472815482081368499$ $3.86\times 10^{-31}$
55 $9.28044435522920809582$ $1.09\times 10^{-31}$ $15$ $23.08868709418032385846$ $1.74\times 10^{-31}$
66 $10.6382769795389180751$ $1.54\times 10^{-31}$ $16$ $24.43655444208668619203$ $2.25\times 10^{-31}$
77 $12.8151786361302353694$ $5.75\times 10^{-32}$ $17$ $25.87097298765072715474$ $1.42\times 10^{-31}$
88 $13.6043079656071037781$ $1.37\times 10^{-31}$ $18$ $26.26140271050477788697$ $1.49\times 10^{-31}$
99 $15.2714728878515316883$ $8.72\times 10^{-32}$ 1919 $27.96504990212936504067$ $2.45\times 10^{-31}$
$10$ $16.1666343336440938522$ $3.22\times 10^{-31}$ 2020 $29.27531093215095949777$ $8.60\times 10^{-32}$
@@
11 $1.86993927318275168341$ $4.52\times 10^{-32}$ $11$ $17.61972287425101875532$ $9.04\times 10^{-32}$
22 $4.70920812840911269662$ $2.64\times 10^{-31}$ $12$ $18.26903309787599031651$ $2.28\times 10^{-31}$
33 $5.81981644686087148921$ $4.85\times 10^{-32}$ $13$ $20.30638125042707154147$ $2.32\times 10^{-31}$
44 $7.38790752248483336704$ $3.58\times 10^{-31}$ $14$ $21.66563972555785204781$ $3.05\times 10^{-31}$
55 $9.44566152389131881370$ $2.60\times 10^{-31}$ $15$ $22.32883476081719799605$ $6.78\times 10^{-32}$
66 $10.7407497975074755000$ $9.60\times 10^{-32}$ $16$ $24.04471027020575246775$ $6.45\times 10^{-32}$
77 $11.9563737277407041057$ $1.59\times 10^{-31}$ $17$ $24.68148321849740900845$ $6.86\times 10^{-33}$
88 $13.4164345305120382816$ $2.22\times 10^{-32}$ $18$ $26.44242345116378413034$ $5.29\times 10^{-31}$
99 $14.8319853387339990685$ $1.08\times 10^{-31}$ 1919 $27.60237248544315232420$ $3.84\times 10^{-31}$
$10$ $16.5192247034936462349$ $2.60\times 10^{-31}$ 2020 $28.77893942445047647073$ $2.36\times 10^{-31}$
@@ Appendix E ARB Certification of Zero Data
All 340 zeros tabulated in Appendices C and B were rigorously certified using the methodology of §A.1A.2. Working precision was 500 bits for the seven auxiliary characters and 1500 bits for $L(s,\chi_{5})$. The following table summarizes the certified upper bounds:
@@
$\chi_{5}$ $\mathbb{Q}(\sqrt{5})$ 200 $6\times 10^{-358}$
$\chi_{2}$ $\mathbb{Q}(\sqrt{2})$ 20 $4\times 10^{-147}$
$\chi_{3}$ $\mathbb{Q}(\sqrt{3})$ 20 $4\times 10^{-147}$
$\chi_{13}$ $\mathbb{Q}(\sqrt{13})$ 20 $4\times 10^{-147}$
$\chi_{6}$ $\mathbb{Q}(\sqrt{6})$ 20 $4\times 10^{-147}$
$\chi_{7}$ $\mathbb{Q}(\sqrt{7})$ 20 $4\times 10^{-147}$
$\chi_{10}$ $\mathbb{Q}(\sqrt{10})$ 20 $4\times 10^{-147}$
$\chi_{11}$ $\mathbb{Q}(\sqrt{11})$ 20 $4\times 10^{-147}$
@@ Appendix F Tables of Zeros of Quadratic Dirichlet $L$-Functions of Small Conductor
Due to its length (approximately 181 pages), this appendix is archived separately on Zenodo:
Appendix F: Tables of Zeros of Quadratic Dirichlet $L$-Functions of Small Conductor.
Available at https://doi.org/10.5281/zenodo.18783098.
The companion document tabulates the first 1000 zeros of $L(s,\chi_{d})$ to 70 decimal places for each of the eight quadratic characters $\chi_{2}$, $\chi_{3}$, $\chi_{5}$, $\chi_{6}$, $\chi_{7}$, $\chi_{10}$, $\chi_{11}$, and $\chi_{13}$ (8000 zeros in total). All zeros were computed using the methodology of §A.2 at 1500-bit working precision and rigorously certified using ARB interval arithmetic (§A.1). Certified upper bounds on $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi_{d})|$ are provided for every zero, with all bounds below $10^{-447}$. This dataset extends and supersedes the tables in Appendices B and C of the present paper.

### 12 Mar 2026 v3

@@ Appendix C Quadratic $L$-Function Zeros to 20 Decimal Places (Seven Discriminants)
11 $4.89997399700703650103$ $6.05\times 10^{-451}$ $11$ $26.95853518038046741968$ $4.22\times 10^{-451}$
22 $7.62842884176939783416$ $4.95\times 10^{-451}$ $12$ $28.09744496063079735461$ $8.97\times 10^{-452}$
33 $10.80658816386171201438$ $9.59\times 10^{-451}$ $13$ $29.93076421015190503722$ $8.58\times 10^{-452}$
44 $12.31054299423652968112$ $7.48\times 10^{-452}$ $14$ $31.63813949132101731704$ $4.01\times 10^{-451}$
55 $15.19575425064512276844$ $5.78\times 10^{-451}$ $15$ $33.84563089515844186412$ $2.62\times 10^{-451}$
66 $17.02228597430834733897$ $3.34\times 10^{-451}$ $16$ $34.74577700617002330947$ $8.15\times 10^{-452}$
77 $18.80595890770714839982$ $6.76\times 10^{-452}$ $17$ $36.54166385177458968906$ $3.99\times 10^{-451}$
88 $21.13164596222134388263$ $1.61\times 10^{-451}$ $18$ $38.77557777387006581423$ $4.34\times 10^{-451}$
99 $23.08384999620054654288$ $9.60\times 10^{-452}$ 1919 $39.78688099766664071794$ $8.55\times 10^{-451}$
$10$ $24.20196355781560161254$ $3.23\times 10^{-451}$ 2020 $41.34220616442193704510$ $9.11\times 10^{-452}$
@@
11 $3.80462763305086509714$ $5.70\times 10^{-451}$ $11$ $25.41163389239269596203$ $5.27\times 10^{-450}$
22 $6.69222332050013115991$ $1.74\times 10^{-450}$ $12$ $27.01394398585067234434$ $6.53\times 10^{-450}$
33 $8.89059295872674150933$ $1.46\times 10^{-450}$ $13$ $28.44220325774560811850$ $6.16\times 10^{-450}$
44 $11.18839274507490915416$ $2.51\times 10^{-450}$ $14$ $30.20400655643888185431$ $7.85\times 10^{-450}$
55 $12.96617880802884537495$ $2.97\times 10^{-450}$ $15$ $33.03713287950650360192$ $8.71\times 10^{-450}$
66 $15.18148087588821673539$ $3.23\times 10^{-450}$ $16$ $35.02737848501270258165$ $5.90\times 10^{-450}$
77 $16.63263327452376212317$ $3.62\times 10^{-450}$ $17$ $35.77804457652351027241$ $8.68\times 10^{-450}$
88 $18.88436945712065080816$ $4.96\times 10^{-450}$ $18$ $37.92681682137521283528$ $4.63\times 10^{-450}$
99 $20.10392819124581857313$ $3.22\times 10^{-450}$ 1919 $38.97399882227494028953$ $6.51\times 10^{-450}$
$10$ $23.56131971313784471634$ $6.15\times 10^{-450}$ 2020 $40.48415475080402043184$ $1.20\times 10^{-449}$
@@
11 $3.11934147900860341390$ $99\times 10^{-450} 11$11 $23.59202788517293715038$ $1.59\times 10^{-449}$
22 $7.23159073941876201502$ $3.75\times 10^{-450}$ $12$ $25.37170440577142621541$ $1.49\times 10^{-449}$
33 $8.62542663503259159734$ $3.76\times 10^{-450}$ $13$ $26.38431352575075899278$ $1.08\times 10^{-449}$
44 $10.33642072623153902983$ $4.67\times 10^{-450}$ $14$ $27.64980842472696812713$ $3.77\times 10^{-449}$
55 $12.61701279102317873036$ $6.95\times 10^{-450}$ $15$ $29.39145603391187839555$ $1.01\times 10^{-449}$
66 $15.14833241700574422460$ $7.72\times 10^{-450}$ $16$ $31.01874209224480391870$ $1.61\times 10^{-449}$
77 $16.27482605749858841789$ $1.03\times 10^{-449}$ $17$ $32.42341728496373366098$ $1.85\times 10^{-449}$
88 $18.75125235623423291645$ $1.01\times 10^{-449}$ $18$ $34.46644233774365394408$ $1.97\times 10^{-449}$
99 $19.54804144334703986992$ $7.62\times 10^{-450}$ 1919 $35.76319303421611596021$ $1.89\times 10^{-449}$
$10$ $20.95918191740503118143$ $1.78\times 10^{-449}$ 2020 $36.72754705153707320059$ $1.52\times 10^{-449}$
@@
11 $2.68865813246759758667$ $04\times 10^{-450} 11$11 $21.55547824218793740139$ $7.57\times 10^{-450}$
22 $5.29243117677719815159$ $2.22\times 10^{-450}$ $12$ $22.66785732205079621070$ $7.90\times 10^{-450}$
33 $9.22463743993556115081$ $2.96\times 10^{-450}$ $13$ $24.40914964223159369359$ $9.44\times 10^{-450}$
44 $10.44572151841292482301$ $3.58\times 10^{-450}$ $14$ $25.67129733212202220561$ $5.62\times 10^{-450}$
55 $12.63115371816651058549$ $4.36\times 10^{-450}$ $15$ $26.75610741237823107923$ $6.79\times 10^{-450}$
66 $13.77715270577349586517$ $3.79\times 10^{-450}$ $16$ $28.86856257427073136901$ $7.97\times 10^{-450}$
77 $15.60365697037025378285$ $5.64\times 10^{-450}$ $17$ $29.21969083370577778290$ $5.98\times 10^{-450}$
88 $17.09898808691940348689$ $7.49\times 10^{-450}$ $18$ $31.08725180031803397862$ $9.52\times 10^{-450}$
99 $18.46329885012140044080$ $7.71\times 10^{-450}$ 1919 $32.42353686755135030793$ $8.63\times 10^{-450}$
$10$ $20.03461157675018829472$ $6.41\times 10^{-450}$ 2020 $33.49155789395814591322$ $1.20\times 10^{-449}$
@@
11 $2.77728324207659233094$ $2\times 10^{-450} 1111$ $20.39328285450433247152$ $7.78\times 10^{-450}$
22 $4.52677408701547610928$ $2.48\times 10^{-450}$ $12$ $21.98889814760568313346$ $8.12\times 10^{-450}$
33 $7.29088506700227937478$ $3.27\times 10^{-450}$ $13$ $24.73886680489813285047$ $9.06\times 10^{-450}$
44 $8.39512407283755124263$ $3.62\times 10^{-450}$ $14$ $26.38179066809542195016$ $9.15\times 10^{-450}$
55 $10.11630685473129779109$ $4.04\times 10^{-450}$ $15$ $27.02957442259405572955$ $1.15\times 10^{-449}$
66 $12.02206203818595380135$ $6.06\times 10^{-450}$ $16$ $28.72073200609003118776$ $8.03\times 10^{-450}$
77 $13.54428637182978741492$ $5.95\times 10^{-450}$ $17$ $30.39449098988413064026$ $1.06\times 10^{-449}$
88 $16.01626147715059280892$ $5.90\times 10^{-450}$ $18$ $31.36586297616383901440$ $1.13\times 10^{-449}$
99 $18.20042527549262445931$ $9.95\times 10^{-450}$ 1919 $32.18417572788333303375$ $1.43\times 10^{-449}$
$10$ $19.53861609989941145162$ $8.33\times 10^{-450}$ 2020 $33.84959183673288697276$ $1.12\times 10^{-449}$
@@
11 $2.48821020930540853564$ $31\times 10^{-450} 11$11 $19.28135808746821972149$ $9.43\times 10^{-450}$
22 $4.03159701418841483370$ $2.20\times 10^{-450}$ $12$ $20.40160148601184344649$ $6.53\times 10^{-450}$
33 $6.33545319091356197519$ $3.46\times 10^{-450}$ $13$ $21.56472815482081368499$ $1.17\times 10^{-449}$
44 $7.98779634495761720535$ $4.61\times 10^{-450}$ $14$ $23.08868709418032385846$ $1.18\times 10^{-449}$
55 $10.63827697953891807504$ $4.03\times 10^{-450}$ $15$ $24.43655444208668619202$ $9.45\times 10^{-450}$
66 $12.81517863613023536936$ $5.16\times 10^{-450}$ $16$ $25.87097298765072715474$ $9.39\times 10^{-450}$
77 $13.60430796560710377807$ $4.11\times 10^{-450}$ $17$ $26.26140271050477788696$ $1.07\times 10^{-449}$
88 $15.27147288785153168831$ $5.54\times 10^{-450}$ $18$ $27.96504990212936504067$ $1.43\times 10^{-449}$
99 $16.16663433364409385216$ $7.44\times 10^{-450}$ 1919 $29.27531093215095949777$ $1.04\times 10^{-449}$
$10$ $18.18218004680212107005$ $1.05\times 10^{-449}$ 2020 $30.53574562468521890003$ $1.38\times 10^{-449}$
@@
11 $1.86993927318275168340$ $2.18\times 10^{-450}$ $11$ $17.61972287425101875532$ $9.14\times 10^{-450}$
22 $4.70920812840911269661$ $3.30\times 10^{-450}$ $12$ $18.26903309787599031650$ $7.78\times 10^{-450}$
33 $5.81981644686087148920$ $2.59\times 10^{-450}$ $13$ $20.30638125042707154146$ $1.11\times 10^{-449}$
44 $7.38790752248483336704$ $4.33\times 10^{-450}$ $14$ $21.66563972555785204780$ $8.68\times 10^{-450}$
55 $9.44566152389131881369$ $5.59\times 10^{-450}$ $15$ $22.32883476081719799605$ $9.52\times 10^{-450}$
66 $10.74074979750747559997$ $4.47\times 10^{-450}$ $16$ $24.04471027020575246774$ $1.10\times 10^{-449}$
77 $11.95637372774070410568$ $5.55\times 10^{-450}$ $17$ $24.68148321849740900845$ $1.19\times 10^{-449}$
88 $13.41643453051203828157$ $7.98\times 10^{-450}$ $18$ $26.44242345116378413033$ $1.35\times 10^{-449}$
99 $14.83198533873399906847$ $1.06\times 10^{-449}$ 1919 $27.60237248544315232420$ $1.44\times 10^{-449}$
$10$ $16.51922470349364623489$ $7.80\times 10^{-450}$ 2020 $28.77893942445047647072$ $1.59\times 10^{-449}$
@@ Appendix E ARB Certification of Zero Data
All 340 zeros tabulated in Appendices C and B were rigorously certified using the methodology of §A.1A.2. Working precision was at least 1500 bits for all characters. The following table summarizes the certified upper bounds:
@@
$\chi_{5}$ $\mathbb{Q}(\sqrt{5})$ 200 $10^{-448}$
$\chi_{2}$ $\mathbb{Q}(\sqrt{2})$ 20 $10^{-451}$
$\chi_{3}$ $\mathbb{Q}(\sqrt{3})$ 20 $10^{-449}$
$\chi_{13}$ $\mathbb{Q}(\sqrt{13})$ 20 $10^{-449}$
$\chi_{6}$ $\mathbb{Q}(\sqrt{6})$ 20 $10^{-449}$
$\chi_{7}$ $\mathbb{Q}(\sqrt{7})$ 20 $10^{-449}$
$\chi_{10}$ $\mathbb{Q}(\sqrt{10})$ 20 $10^{-449}$
$\chi_{11}$ $\mathbb{Q}(\sqrt{11})$ 20 $10^{-449}$
@@ Appendix F Tables of Zeros of Quadratic Dirichlet $L$-Functions of Small Conductor
Due to its length (approximately 186 pages), this appendix is archived separately on Zenodo and GitHub: Appendix F: Tables of Zeros of Quadratic Dirichlet $L$-Functions of Small Conductor.
The companion document tabulates between 1004 and 1044 zeros of $L(s,\chi_{d})$ to 70 decimal places for each of the eight quadratic characters $\chi_{2}$, $\chi_{3}$, $\chi_{5}$, $\chi_{6}$, $\chi_{7}$, $\chi_{10}$, $\chi_{11}$, and $\chi_{13}$ (8244 zeros in total). All zeros were computed using the methodology of §A.2 at a minimum of 1500-bit working precision and rigorously certified using ARB interval arithmetic (§A.1). Certified upper bounds on $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi_{d})|$ are provided for every zero, with all bounds below $10^{-409}$. This dataset extends and supersedes the tables in Appendices B and C of the present paper.